OpenStudy (anonymous):

Gwen has 3 red erasers, 4 green erasers, 2 black erasers, 5 purple erasers, and 6 yellow erasers in her book bag. What is the probability Gwen will reach in her book bag and select a purple eraser?

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

would it not be 1/4

OpenStudy (cwrw238):

there are 20 erasers altogether of which 5 are purple so probability of picking a purple = 5/20 = 1/4

OpenStudy (anonymous):

thank u i wasint to sure cwrw238